jhonyy9 (jhonyy9):

every primes graters than 2 can writing in the form (2 on exponent (n+1) +1). True or False ?

OpenStudy (anonymous):

\[2^{n+1}+1\]?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

so prove it than ...

jhonyy9 (jhonyy9):

n>=0

OpenStudy (anonymous):

It's false i think

OpenStudy (anonymous):

for example, it does't work for 7

OpenStudy (anonymous):

If n is supposed to be an integer, 7 would be a fine counterexample.

OpenStudy (anonymous):

here is my proof. 109 is prime

jhonyy9 (jhonyy9):

moneybird you can prove it ?

OpenStudy (anonymous):

actually i like moneybirds proof better. only need to go up to 7

jhonyy9 (jhonyy9):

so than the primes not can never writing in the form 2 on exponent (n) +1 or ...how ?

OpenStudy (anonymous):

5 can be written that way as \[2^2+1\]

jhonyy9 (jhonyy9):

just in the form (2n+1) ?

OpenStudy (anonymous):

The statement says every. If you have a counterexample, then the statement is false.

jhonyy9 (jhonyy9):

and 7=2*2+2+1

OpenStudy (anonymous):

i think there is some confusion in this question . why would anyone write \[2^{n+1}+1\] instead of \[2^n+1\]?

jhonyy9 (jhonyy9):

not is just for we can getting when n=0 so yes sorry i have wrote firstly 2 on exponent n-1 +1

jhonyy9 (jhonyy9):

when n=1 than we get 2 on exponent 0 =1 so than will be 1+1=2

jhonyy9 (jhonyy9):

than 2+2=4 and hence in this way we can getting 4 like sum of two primes

OpenStudy (anonymous):

i am still confused by the question. it is not true that for all n \[2^n+1\] is prime and it is not true that for all primes \[p=2^n+1\]

OpenStudy (anonymous):

for example \[2^3+1=9\] not prime and \[7\neq 2^n+1\]

jhonyy9 (jhonyy9):

thank you satellite73 so than 2 on exponent n + 3 will be right always ?

jhonyy9 (jhonyy9):

i think yes